find the mean median modes and range for the following data sets

7,10,11,12,11,11,13,8,22,23
Answer #1

Ans:

Sorted data:

7,8,10,11,11,11,12,13,22,23

sum=128

n=10

mean=128/10=12.8

median=11

mode=11

Range=23-7=16